OpenStudy (anonymous):

find the tangent line to x^2 + 2y = 8 and parallel to 2x - y = 4

OpenStudy (anonymous):

need to use calculus

OpenStudy (anonymous):

just give me the steps and i'll try solving for it.. thanks

OpenStudy (phi):

find the derivative of the curve, that will give you an equation for the slope at any point on the curve. find the slope of the line 2x - y - 4 use that slope in the first equation

OpenStudy (anonymous):

will i equate the curve to y first? then defferentiate it?

OpenStudy (phi):

yes, solve for y (because you want to find dy/dx = slope) or you can use implicit differentiation if you know how.

OpenStudy (anonymous):

i arrive at two different answers.. 1st what i did is 2y = 8 - x^2 .. then i devide 2.. so what i got is y = (8 - x^2)/2 .. derive.. i got dy/dx = -x. then i tried doing implicit i got dy/dx = -2x so which is correct?

OpenStudy (phi):

implict: x^2 + 2y = 8 2 x dx/dx + 2 dy/dx = 0 2x + 2 dy/dx =0 x + dy/dx =0 dy/dx = -x

OpenStudy (anonymous):

owh.. yes.. sorry.. got confused

OpenStudy (phi):

Here is a plot

OpenStudy (anonymous):

what is the purpose of dy/dx = -x ?

OpenStudy (phi):

that is the slope of the curve (as a function of x) at x=0 the slope is -0 = 0 you want to find the x on the curve where the slope matches the line they gave you

OpenStudy (anonymous):

ok.. then solve for the slope of 2x - y - 4 = 0.. i got 2.. what next?

OpenStudy (phi):

find the tangent line to x^2 + 2y = 8 where on the curve x^2 +2y =8 does it have a slope of 2 ?

OpenStudy (anonymous):

got confused @.@

OpenStudy (phi):

when you get a chance, watch this http://www.khanacademy.org/math/calculus/differential-calculus/derivative_intro/v/calculus--derivatives-1--new-hd-version meanwhile, dy/dx = -x (the derivative of x^2 + 2y=8 ) tells you the slope of the tangent line of that curve at any point x

OpenStudy (phi):

you want the tangent line that has a slope of 2 (to match the slope of the line 2x - y = 4) m= -x 2 = -x so x= -2 at x=-2 the tangent line to the curve will have a slope of 2 (see post up above of the plot)

OpenStudy (phi):

now you need to find the y value on the curve when x= -2 that will give you a point on the curve (-2, ?) you have the slope m=2 now you can write down the equation of the line through that point with slope =2

OpenStudy (anonymous):

where will i substitute the value -2?

OpenStudy (phi):

if you want the (x,y) values of a point on the curve, and you have x= -2, how do you find y ?

OpenStudy (anonymous):

substitute -2 from x.. 2^2 + 2y = 8 ... 2y = 4 .. y = 2.. is that correct?

OpenStudy (phi):

yes

OpenStudy (anonymous):

then i will just need to use point-slope formula to find the equation of the line perpendicular to 2x - y - 4 = 0 ?

OpenStudy (phi):

though you should write it as (-2)^2 + 2y = 8

OpenStudy (phi):

find the tangent line to x^2 + 2y = 8 and *parallel* to 2x - y = 4

OpenStudy (anonymous):

the tangent line will have the pass the point (-2,4) ?

OpenStudy (phi):

? I thought you found y=2 at x=-2

OpenStudy (anonymous):

yes.. sorry.. got wrong with my typing.. (-2,2) so y-y1=m(x-x1) y - 2 = 2(x+2) y = 2x + 6.. that is the equation of the tanget line

OpenStudy (phi):

looks good

OpenStudy (anonymous):

is that correct?? so that is perpendicular to 2x-y-4 because they have the same slope..

OpenStudy (phi):

you mean parallel perpendicular means they form a right angle (and the slopes are negative reciprocals)

OpenStudy (anonymous):

ahh yes.. but is my answer correct already? so i can ask some questions again

OpenStudy (phi):

yes you found the tangent line that is parallel to the given line

OpenStudy (anonymous):

thanks ! you're a great help, and a good tutor..